OpenStudy (anonymous):

3 island in the hawaiian island chain are hawaii , maui, and kauai. together they total 5300 square miles. hawaii is 3293 sq. miles bigger than maui and maui is 177 sq. miles larger than kauai . whats the area of each island?

OpenStudy (shadowfiend):

So you have three islands with areas h (hawaii), m (maui), and k (kauai). h + m + k = 5300 m + 3293 = h m = k + 177 Can you solve that?